the length of a rectangle is 3 more than twice its width. if the perimeter is 54 inches, what is the width of the rectangle?
Answer by algebrahouse.com(1659)

"the length of a rectangle is 3 more than twice its width. if the perimeter is 54 inches, what is the width of the rectangle?"

x = width
2x + 3 = length {length is 3 more than twice width}

Perimeter of a rectangle is 2(width) + 2(length)

2x + 2(2x + 3) = 54 {perimeter = 2(width) + 2(length)}
2x + 4x + 6 = 54 {used distributive property}
6x + 6 = 54 {combined like terms}
6x = 48 {subtracted 6 from both sides}
x = 8 {divided both sides by 6}
2x + 3 = 19 {substituted 8, in for x, into 2x + 3}

width = 8 and length = 19

For more help from me, visit: www.algebrahouse.com